OpenStudy (anonymous):

Solve the polynomial inequality. Express the solution set in interval notation: (-x - 4)/(x + 3) ≤ 0

OpenStudy (anonymous):

Any ideas?

OpenStudy (anonymous):

You should try drawing the graph of this function. Plot the asymptote(s) first and then from there consider whether your function will be negative (i.e. <= 0) to the left or right of your asymptote(s).

OpenStudy (anonymous):

Is there really no other, simpler approach?

OpenStudy (anonymous):

find critical points... i.e. points that result in zero or undefined. \[\frac{-x-4}{x+3}=0\]-x-4 gives x=-4 x+3=0 gives x doesn't equal -3